OpenStudy (anonymous):

use log differentiation : y=(lnx)^(1/x)

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

start with \[\frac{1}{x}\ln(\ln(x))\] then take the derivative using the product and chain rule

OpenStudy (misty1212):

when you are done, multply your answer by \[\ln(x)^{\frac{1}{x}}\]